Solve (D^2 -4D+3)y = e^x cos 2x + cos3x 


1
Expert's answer
2021-05-25T18:07:54-0400

We have given the differential equation,


"(D^2 -4D+3)y = e^x cos 2x + cos3x"


Auxiliary equation can be given as,

"m^2-4m+3 =0\\

m^2-3m-m+3 = 0\\

m(m-3)-1(m-3) = 0\\

m=1,2"  

Hence, "CF = C_1e^x+C_2e^{2x}"

Particular integral can be calculated as,


"PI = \dfrac{cos3x}{-9-4D+3}+ \dfrac{e^xcos2x}{(D+1)^2-4(D+1)+3}"


We know,

"\dfrac{1}{f(D^2)}sinax = \dfrac{sinax}{f(-a^2)}" and "\dfrac{1}{f(D^2)}cosax = \dfrac{cosax}{f(-a^2)}"


Now Multiplying by "(D-3)" in numerator and denominator of first part of particular integral.


"PI = \dfrac{cos3x(D-3)}{-2(D^2-9)}+ \dfrac{e^xcos2x}{D^2-2D}"


After solving we get,



"PI = \dfrac{-3sin3x-3cos3x}{36} + \dfrac{e^x(-2sin2x-2cos2x)}{16}"


Hence , the complete solution is

"y = C_1e^x+C_2e^{2x}+ \dfrac{-3sin3x-3cos3x}{36} + \dfrac{e^x(-2sin2x-2cos2x)}{16}"